# Accelerating Autoregressive Speech Synthesis Inference With Speech Speculative Decoding

### v1

@@ 1 Introduction
Inspired by speculative decoding techniques originally developed for text-based LLMs [22, 23], we present a novel approach of Speech Speculative Decoding (SSD) to accelerate autoregressive speech synthesis. Speculative decoding leverages a small draft model to generate candidate token sequences autoregressively, which are then verified by the larger target model in parallel. This approach significantly reduces inference latency while preserving synthesis quality [24]. Unlike text tokens, which often have clear, one-to-one mappings with words or phonemes, speech tokens exhibit diverse and variable mappings, as audio segments that sound nearly identical can have different token-level representations. This variability stems from the continuous and dynamic nature of speech signals. Traditional speculative decoding [22, 23], designed for text, may result in suboptimal efficiency for speech synthesis, as their acceptance criteria fail to account for these nuanced relationships. To address this, we propose an improved acceptance criterion tailored to speech tokens, enabling more efficient verification of candidate sequences generated by the draft model. The adaptation in SSD finally achieves a significant speedup of 1.4x compared with conventional autoregressive decoding, while maintaining high fidelity and naturalness.
@@ 2 Methodology
In this section, we demonstrate the proposed SSD acceleration framework, which comprises two core components: a target model (the large model requiring acceleration) and a draft model (a lightweight auxiliary model). As illustrated in Figure 2, the process unfolds in three key phases: 1) Autoregressively generating $m$ candidate tokens using the draft model. 2) Parallel verification of these tokens in the target model via SSD. 3) Applying a speculative decoding-based acceptance strategy to validate the token sequences.
@@ 2.1 Autoregressive Language Model for Speech Synthesis
In autoregressive speech synthesis system, a language model is applied to predict the discrete tokens of speech. Our approach builds upon the open-source Cosyvoice framework [4, 5] for autoregressive speech synthesis, which uses language model to predict single stream of discrete speech tokens and then uses flow-matching model as the speech decoder.
Given a text-speech pair $\{T,A\}$, where $T$ denotes text sequence and $A$ represents the speech signal, text $T$ is tokenized into a more fine-grained sequence $T=\{t_{1},t_{2},\ldots,t_{N}\}$ denotes text squence of length $N$ via some methods such as grapheme-to-phoneme (g2p) transformation or BPE-based text tokenizer and the speech encoder discretizes the continuous speech signal into a sequence of discrete tokens $Y=\{y_{1},y_{2},\ldots y_{L}\}$ of length $L$. Then the language model predicts $Y$ autoregressively taken text $T$ as input. To input all the information into the language model, text tokens and speech tokens are typically concatenated into a single sequence via some special tokens and then uniformly fed into the model.
@@
where $Y_{pt}$ denotes the audio prompt, $T_{pt}$ represents the transcription of audio prompt, and $T_{tgt}$ means the target text. Traditional autoregressive models suffer from speed limitations due to their token-by-token serial generation process. This sequential dependency significantly reduces computational efficiency and brings to high latency.
Input: Input prompt $T_{in}$, target model $M_{q}$, draft model $M_{p}$, draft length $L_{d}$, target length $L_{t}$, tolerance $\beta$
@@ 2.2 Speech Speculative Decoding
As shown in Algorithm 1, Speculative decoding accelerates autoregressive generation by leveraging a draft model $M_{p}$ to propose candidate tokens and a target model $M_{q}$ to verify them. In speech synthesis, we adapt this framework to address the inherent ambiguity of speech tokens, which lack explicit alignment with linguistic units. Our key innovation lies in introducing a tolerance mechanism to enhance the flexibility of decoding while preserving the quality of the output.
@@
Unlike text generation where tokens have discrete semantics and one-to-one mappings with words or phonemes, speech tokens exhibit continuous and overlapping characteristics. Strict acceptance criteria in standard speculative decoding often lead to unnecessary rejections, as minor distribution discrepancies may not perceptually affect speech quality. To address this, we introduce a tolerance factor $\beta\geq 0$ that relaxes the acceptance threshold:
@@
When rejection occurs at position $j$, we resample from the adjusted distribution:
$p_{new}=\text{normlize}(\max(0,q_{j}^{s}-p_{j}^{d})),$ (6)
@@ 2.3 Lightweight Draft Model Construction
Modern autoregressive speech synthesis systems rely on a single large target model, posing a key challenge for speculative decoding: constructing a lightweight draft model that aligns with the vocabulary of target model. To address this, we design the draft model as a shallower variant of the target architecture, sharing its core structure but with fewer Transformer layers. Crucially, the draft model is initialized using the pre-trained upper layers of the target model (e.g., layers 18–24 in a 24-layer setup), enabling immediate compatibility with its linguistic representations while avoiding cold-start training.
@@
where $U_{tgt}=\{u_{1},...,u_{L_{tgt}}\}$ is the target sequence, $u_{<l}=\{u_{1},...,u_{l-1}\}$ and $L_{tgt}$ is the length of the target sequence.
@@ 3 Experiments
In this section, we will introduce the experimental settings, including the implementation details, dataset, and metrics.
@@ 3.1 Implementation Details
We use CosyVoice 2 [5] as the target model, which is a highly efficient TTS model consisting of 24 layers of Transformer adapted from Qwen2.5 (0.5B) [27]. Based on this model, we design a draft model consisting of 8 layers of Transformer. We follow the setting of CosyVoice 2. Each Transformer layer typically has a multi-head attention mechanism, where the number of heads is set to 14. The feed-forward layer in each Transformer layer has a hidden dimension of 896. We train the draft model on LibriTTS [28] while keeping the parameters of CosyVoice 2 frozen. To accelerate the training process, we load the pre-trained parameters of CosyVoice 2 to initialize the model. For the Transformer layers, we utilize the parameters of the first two layers and the last six layers of CosyVoice 2. During the training of the draft model, we only train the first two layers of the Transformer and the classification layer. The length of draft tokens as $L_{d}$ is set to 3 and the tolerance factor $\beta$ is set to 0.4. We conduct all the experiments on NVIDIA A100-SXM4-40GB GPU.
@@ 3.2 Dataset
Following [29], we evaluate SSD on the subset of LibriTTS. This testset, which covers 37 speakers, contains 500 utterances selected from the ”test-clean” subset of LibriTTS. Its purpose is to evaluate the zero-shot adaptation ability for entirely new and unseen speakers. Moreover, for each speaker in the testset, there is a short voice prompt lasting approximately 3 seconds.
@@ 3.3 Evaluation Metrics
For subjective evaluation, we conduct two MOS (Mean Opinion Score) experiment. First, we focus on the similarity (SIM-MOS), which provides insights into how closely the synthetic voice mimics the unique vocal characteristics of the target speaker. Then, we evaluate naturalness of the speech (NAT-MOS), meaning how close the synthesized voice is to a real-human voice and detecting any signs of artificial synthesis.
@@ 4 Results and Analysis
We will present the subjective and objective experimental results of the audio in detail, and conduct an exhaustive analysis of SSD. We apply SSD to CosyVoice 2 and compare the result with draft model and CosyVoice 2 using autoregressive decoding to synthesize speech.
@@ 4.2 Subjective Evaluation
The subjective evaluation results in Table 2 demonstrate that SSD achieves comparable performance to the original CosyVoice 2 using autoregressive decoding in terms of both similarity (SIM-MOS) and naturalness (NAT-MOS). While the draft model alone achieves the lowest scores, highlighting its limitations in independent high-quality speech generation, SSD bridges this gap by leveraging the target model for verification. These results confirm that SSD effectively preserves the synthesis quality of the target model while accelerating inference, validating it as a practical solution for efficient autoregressive speech synthesis without compromising perceptual quality.

### v2

@@ 1 Introduction
Speculative decoding [22, 23] accelerates text-based large language models (LLMs) by employing a small draft model to autoregressively generate candidate token sequences, which are then verified in parallel by a larger target model. Inspired by these approaches, there have been notable attempts [24, 25] to apply speculative decoding to accelerate autoregressive speech synthesis. However, they integrate multiple draft heads directly into the target model and tend to require fine-tuning the target model itself to enhance draft head quality. In contrast, we introduce Speech Speculative Decoding (SSD) with an independent draft architecture that maintains high-quality predictions without requiring target model fine-tuning. Moreover, unlike text tokens, which often have clear one-to-one mappings with words or phonemes, speech tokens exhibit diverse and variable mappings, as audio segments that sound nearly identical can have different token-level representations. This variability stems from the continuous and dynamic nature of speech signals. Traditional speculative decoding [22, 23], designed for text, may result in suboptimal efficiency for speech synthesis, as their acceptance criteria fail to account for these nuanced relationships. To address this, we propose an improved acceptance criterion tailored to speech tokens, enabling more efficient verification of candidate sequences generated by the draft model. The adaptation in SSD finally achieves a significant speedup of 1.4$\times$ compared to conventional autoregressive decoding, while maintaining high fidelity and naturalness.
@@ 2 Methodology
In this section, we demonstrate the proposed SSD acceleration framework, which comprises two core components: a target model (the large model requiring acceleration) and a draft model (a lightweight auxiliary model). As illustrated in Figure 2, the process unfolds in three phases: 1) Autoregressively generating $m$ candidate tokens using the draft model, 2) Parallel verification of these tokens in the target model via SSD, 3) Applying a speculative decoding-based acceptance strategy to validate the token sequences.
@@ 2.1 Autoregressive Language Model for Speech Synthesis
In an autoregressive speech synthesis system, a language model is applied to predict discrete tokens of speech. Our approach builds upon the open-source CosyVoice framework [4, 5] for autoregressive speech synthesis, which uses language model to predict single stream of discrete speech tokens and then uses flow-matching model as the speech decoder.
Given a text-speech pair $\{T,A\}$, where $T$ denotes text sequence and $A$ represents the speech signal, text $T$ is tokenized into a more fine-grained sequence $T=\{t_{1},t_{2},\ldots,t_{N}\}$ of length $N$ via some methods such as grapheme-to-phoneme (g2p) transformation or BPE-based text tokenizer, and the speech encoder discretizes the continuous speech signal $A$ into a sequence of discrete tokens $Y=\{y_{1},y_{2},\ldots y_{L}\}$ of length $L$. Then the language model predicts $Y$ autoregressively taken text $T$ as input. To input all the information into the language model, text tokens and speech tokens are typically concatenated into a single sequence via some special tokens and then uniformly fed into the model.
@@
where $Y_{pt}$ denotes the audio prompt, $T_{pt}$ represents the transcription of the audio prompt, and $T_{tgt}$ means the target text. Traditional autoregressive models suffer from speed limitations due to their token-by-token serial generation process. This sequential dependency significantly reduces computational efficiency and brings to high latency.
Input: Prompt $T_{in}$, target model $M_{q}$, draft model $M_{p}$, draft length $L_{d}$, target length $L_{t}$, tolerance factor $\beta$
@@ 2.2 Speech Speculative Decoding
As shown in Algorithm 1, speculative decoding accelerates autoregressive generation by leveraging a draft model $M_{p}$ to propose candidate tokens and a target model $M_{q}$ to verify them. In speech synthesis, we adapt this framework to address the inherent ambiguity of speech tokens, which lack explicit alignment with linguistic units. To enhance the flexibility of speculative decoding for speech synthesis without compromising output quality, we introduce a tolerance factor into the criteria.
@@
Unlike text generation where tokens have discrete semantics and one-to-one mappings with words or phonemes, speech tokens exhibit continuous and overlapping characteristics. Strict acceptance criteria in standard speculative decoding often lead to unnecessary rejections, however minor distribution discrepancies may not perceptually affect speech quality. Therefore, we introduce a tolerance factor $\beta\geq 0$ that relaxes the acceptance threshold:
@@
When rejection occurs at position $i$, we resample token $x^{\prime}_{i}$ from the adjusted distribution:
$\text{Resample }x^{\prime}_{i}\sim p_{new}=\text{normlize}(\max(0,q_{i}^{s}-p_{i}^{d})),$ (6)
@@ 2.3 Lightweight Draft Model Construction
Modern autoregressive speech synthesis systems rely on a single large target model, posing a challenge for speculative decoding: constructing a lightweight draft model that aligns with the vocabulary of target model. To address this, we design the draft model as a shallower variant of the target architecture, sharing its core structure but with fewer Transformer layers. Crucially, the draft model is initialized using the pre-trained upper layers of the target model (e.g., layers 19–24 in a 24-layer setup), enabling immediate compatibility with its linguistic representations while avoiding cold-start training.
@@
where $U_{tgt}=\{u_{1},...,u_{L_{tgt}}\}$ is the target sequence, $u_{<l}=\{u_{1},...,u_{l-1}\}$, and $L_{tgt}$ is the length of the target sequence.
@@ 3.1 Implementation Details
We use CosyVoice 2 [5] as the target model, which is a highly efficient TTS model consisting of 24 layers of Transformer adapted from Qwen2.5 (0.5B) [28]. Based on this model, we design a draft model consisting of 8 layers of Transformer. We follow the setting of CosyVoice 2: each Transformer layer typically employs a multi-head attention mechanism with the number of heads set to 14, and the feed-forward layer in each Transformer layer has a hidden dimension of 896. We train the draft model on LibriTTS [29] while keeping the parameters of CosyVoice 2 frozen. To accelerate the training process, we load the pre-trained parameters of CosyVoice 2 to initialize the model. For the Transformer layers, we utilize the parameters of the lower two layers and the upper six layers of CosyVoice 2. During the training of the draft model, we only train the first two layers of the Transformer and the classification layer. The length of draft tokens as $L_{d}$ is set to 3 and the tolerance factor $\beta$ is set to 0.4. We conduct all the experiments on NVIDIA A100-SXM4-40GB GPU.
@@ 3.2 Dataset
Following [30], we evaluate SSD on the subset of LibriTTS. This testset, which covers 37 speakers, contains 500 utterances selected from the “test-clean” subset of LibriTTS. Its purpose is to evaluate the zero-shot adaptation ability for entirely new and unseen speakers. Moreover, for each speaker in the testset, there is a short voice prompt lasting approximately 3 seconds.
@@ 3.3 Evaluation Metrics
For subjective evaluation, we conduct two Mean Opinion Score (MOS) tests on a 5-point scale with 1-point intervals (1: lowest, 5: highest), involving 30 raters each evaluating 10 randomized samples. First, we focus on the similarity (SIM-MOS), which provides insights into how closely the synthetic voice mimics the unique vocal characteristics of the target speaker. Then, we evaluate naturalness of the speech (NAT-MOS), meaning how close the synthesized voice is to a real-human voice and detecting any signs of artificial synthesis.
@@ 4 Results and Analysis
We apply SSD to CosyVoice 2 and compare the result with draft model and CosyVoice 2 using autoregressive decoding to synthesize speech.
@@ 4.2 Subjective Evaluation
Results in Table 2 demonstrate that SSD achieves comparable performance to the original CosyVoice 2 using autoregressive decoding in terms of both similarity (SIM-MOS) and naturalness (NAT-MOS). While the draft model alone achieves the lowest scores, highlighting its limitations in independent high-quality speech generation, SSD bridges this gap by leveraging the target model for verification. These results confirm that SSD effectively preserves the synthesis quality of the target model while accelerating inference, validating it as a practical solution for efficient autoregressive speech synthesis without compromising perceptual quality.
@@ 5 Conclusions
6 Acknowledgements
This work is supported by National Natural Science Foundation of China (62076144), National Social Science Foundation of China (13&ZD189) and Shenzhen Science and Technology Program (JCYJ20220818101014030).